# A Comprehensive Survey with Critical Analysis for Deepfake Speech Detection

### v1

@@ VI-A2 Future directions
Multilingual and Multimodal Datasets: To address the issue of language diversity, future datasets should include a broader range of languages, accents, and dialects. This variety will enable detection models to better handle diverse linguistic and phonetic features across different languages, ensuring their stability in multilingual contexts and their effectiveness in developing global solutions.
Moreover, deepfake content in real-world scenarios often includes both audio and video elements, rather than just audio. Therefore, integrating multimodal datasets that combine both audio and video deepfakes is a crucial direction for future research. This integration enhances detection capabilities by allowing models to identify anomalies across multiple data types, improving their effectiveness in combating increasingly sophisticated forgeries
@@ VI-F2 Future directions:
VI-G The availability of Deepfake Speech Detection tools

### v4

@@ VI-A2 Future directions
Multilingual and Multimodal Datasets: To address the issue of language diversity, future datasets should include a broader range of languages, accents, and dialects. This variety will enable detection models to better handle diverse linguistic and phonetic features across different languages, ensuring their stability in multilingual contexts and their effectiveness in developing global solutions. Moreover, deepfake content in real-world scenarios often includes both audio and video elements, rather than just audio. Therefore, integrating multimodal datasets that combine both audio and video deepfakes is a crucial direction for future research. This integration enhances detection capabilities by allowing models to identify anomalies across multiple data types, improving their effectiveness in combating increasingly sophisticated forgeries
@@ VI-F2 Future directions
VI-G Feature-free deepfake detection
@@ VI-G1 Open challenges
Deepfake detection faces the usual challenge of the cat-mouse logic of an attack-defense arms race, which is due to the fact that as soon as a feature is identified for detection, it can as quickly be neutralized in the next generation synthesis models. The only way to break this cycle is to develop feature-free detection approaches.
VI-G2 Future directions
For example, Bloom (bloomsocialanalytics.com) proposed a feature-free approach that uses the very same synthesis technologies used to produce deepfakes for their own detection. The idea is based on the intuition that an AI model can reproduce speech produced by an AI more easily than by a human, because reality is always more complex than its model. In other words, real speech contains chaotic components that won’t be perfectly captured by AI models. The proposed method consists of the training and detection phases. The training phase uses an advanced neural voice cloning system to synthesize voice samples based on the target speech files whose authenticity needs to be verified, and then computes a similarity metric between the target speech (authentic or synthetic) and the cloned speech. This distance distribution is used to find the optimal classification threshold, which is then applied to compute the likelihood of authenticity during the detection phase.
VI-H The availability of Deepfake Speech Detection tools
VI-H1 Open challenges
@@ VII Conclusion
ACKNOWLEDGMENTS
The work described in this paper is performed in the H2020 project STARLIGHT (“Sustainable Autonomy and Resilience for LEAs using AI against High priority Threats”). This project has received funding from the European Union’s Horizon 2020 research and innovation program under grant agreement No 101021797.
The work in this paper has further received funding from the European Union - European Defence Fund under GA no. 101121418 (EUCINF). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them.